### v1

# Fuzzy Implicative Rules: A Unified Approach

Abstract
Rule mining algorithms are one of the fundamental techniques in data mining for disclosing significant patterns in terms of linguistic rules expressed in natural language. In this paper, we revisit the concept of fuzzy implicative rule to provide a solid theoretical framework for any fuzzy rule mining algorithm interested in capturing patterns in terms of logical conditionals rather than the co-occurrence of antecedent and consequent. In particular, we study which properties should satisfy the fuzzy operators to ensure a coherent behavior of different quality measures. As a consequence of this study, we introduce a new property of fuzzy implication functions related to a monotone behavior of the generalized modus ponens for which we provide different valid solutions. Also, we prove that our modeling generalizes others if an adequate choice of the fuzzy implication function is made, so it can be seen as an unifying framework. Further, we provide an open-source implementation in Python for mining fuzzy implicative associative rules. We test the applicability and relevance of our framework for different real datasets and fuzzy operators.
Index Terms:
Knowledge discovery in databases (KDD), Rule-based models, Implicative rules, Association rules, Fuzzy logic, Fuzzy implication function.
I Introduction
Data mining or Knowledge Discovery in Databases (KDD) is defined as the automatic extraction of patterns representing knowledge implicitly stored or captured in data [1]. Up to now, a wide variety of data mining techniques have been introduced and developed [2]. These techniques are usually divided into two types: exploratory and predictive. Exploratory data analysis focuses on searching relations between objects of a dataset (clustering, association rules, linguistic summaries…) whereas predictive data analysis aims to extract knowledge from discovered data with the intent to predict or classify unknown examples (classification, regression…). Within this context, rule-based algorithms have been one of the top choices for knowledge extraction because of their human-understandable output. Furthermore, with the emergence research area of eXplainable Artificial Intelligence (XAI) [3] they have become even more popular since they are highly interpretable models which are an appealing solution to provide an easily understandable representation of complex black-box models [4, 5].
In rule-based models, the output is given in terms of rules which are usually represented as $A\Rightarrow C$ in which $A$ is called the antecedent and $C$ is called the consequent. Although in these techniques the output has the same representation, depending on the knowledge to be captured we can find several rule mining techniques, just to mention some of them: (i) association rules: it is interested in capturing situations in which if $A$ is satisfied, then $C$ is likely to occur also [6]; (ii) classification rules: it aims to obtain a set of rules which effectively classify items into predefined classes [7]; (iii) sequential rules: it searches for rules that indicate that if some event(s) occurred, some other event(s) are likely to occur [8]; (iv) subgroup discovery: it aims to obtain rules with the most unusual statistical characteristics with respect to a property of interest [9]
[…]
Further, it is well known that the theoretical development of fuzzy operators has skyrocketed in the last two decades. Indeed, the introduction of new fuzzy operators and the study of their properties is nowadays so vast that there exist many different families from among which we can choose. For instance, only in the case of fuzzy implication functions it has been recently gathered that more than 100 families have been introduced in the literature [12]. This fact, although clearly positive as it indicates great flexibility, also carries with it the intrinsic problem of deciding which operator is the most suitable for each application. In this sense, Zimmerman defines eight different criteria for an adequate selection of the involved aggregation operators in a specific model [13], although authors like Mendel point out that these criteria are rather subjective to be successfully implemented in engineering applications [14]. Thus, the adequate selection of the set of operators is still a hot topic of discussion nowadays.
In this paper, we revisit the definition given in [11] for modeling fuzzy implicative rules and we contextualize it with respect to the advances in fuzzy operators made in recent years. Also, we generalize the four most used quality measures (coverage, support, confidence and weighted relative accuracy) to this new setting. Then, we study which properties should the pair of operators $(T,I)$ satisfy to obtain a coherent behavior of these measures. One of the key points of this study is to prove that for the support to be monotone with respect to the refinements of a rule we have to introduce an additional property of fuzzy implication functions related to the generalized modus ponens. Since up to our knowledge this property had not been previously introduced in the literature, we have studied it in the paper and we have proposed a list of families of operators that can be considered in this framework.
Further, we provide an open-source implementation in Python to mine fuzzy implicative associative rules according to the proposed framework. The implementation allows a custom selection of the fuzzy operators and fuzzy partitions. Finally, we have tested our algorithm in different publicly available databases and admissible pairs of operators. From the results, we have exposed that our perspective provides valuable knowledge which is different to that obtained by other approaches.
The structure of the paper is as follows. First, in Section II we include basic results and definitions. In Section III we introduce fuzzy rules modeled as conditionals, we define different quality measures and we study which properties should the fuzzy operators satisfy. In Section IV we provide solutions for the joint restrictions on the fuzzy operators. In Section V the experimental results are exposed. The paper ends in Section VI with some conclusions and future work.
II Preliminaries
In this section, we provide some basic definitions and results related to fuzzy operators that are used throughout this paper. However, we assume that the reader is familiar with basic concepts of fuzzy sets, fuzzy linguistic variables, fuzzy partitions and fuzzy operators (for further information about these topics, the reader can consult [15, 16, 17]).
Fuzzy conjunctions and disjunctions are defined as increasing binary functions $C,D:[0,1]^{2}\to[0,1]$ such that $C(0,1)=C(1,0)=0$ and $D(0,1)=D(1,0)=1$, respectively. However, it is common to consider commutative, associative operators with neutral element which are called t-norms and t-conorms.
Definition 1 ([16]).
A t-norm (resp. t-conorm) is a binary function which is commutative, associative, increasing in both variables and 1 is its neutral element (resp. 0 is its neutral element).
Example 2.
The following binary functions are t-norms:
  • minimum: $T_{\bm{M}}(x,y)=\min\{x,y\}$.
  • algebraic product: $T_{\bm{P}}(x,y)=x\cdot y$.
  • Łukasiewicz: $T_{\bm{LK}}(x,y)=\max\{x+y-1,0\}$.
Further, fuzzy implication functions are the generalization of the logical conditional to fuzzy logic.
Definition 3 ([17]).
A binary operator $I:[0,1]^{2}\to[0,1]$ is said to be a fuzzy implication function if it satisfies:
$I(x,z)\geq I(y,z)$ when $\ x\leq y$, for all $z\in[0,1]$.
$I(x,y)\leq I(x,z)$ when $\ y\leq z$, for all $x\in[0,1]$.
$I(0,0)=I(1,1)=1$ and $\ I(1,0)=0.$
From Definition 3 it is straightforward to see that if $I$ is a fuzzy implication function then $I(0,x)=I(x,1)=1$ for all $x\in[0,1]$. However, the sections $I(\cdot,0)$ and $I(1,\cdot)$ are not fixed by the definition. In fact, $I(\cdot,0)$ induces a fuzzy negation $N_{I}$ called the natural negation of $I$.
Example 4.
The following binary functions are fuzzy implication functions:
  • Łukasiewicz: $I_{\bm{LK}}(x,y)=\min\{1,1-x+y\}$.
  • Goguen: $I_{\bm{GG}}(x,y)=\left\{\begin{array}[]{ll}1&\text{ if }x\leq y,\\ \frac{y}{x}&\text{ if }x>y.\end{array}\right.$
  • Gödel: $I_{\bm{GD}}(x,y)=\left\{\begin{array}[]{ll}1&\text{ if }x\leq y,\\ y&\text{ if }x>y.\end{array}\right.$
Usually, additional conditions on fuzzy implication functions are considered. Among those in the literature which are relevant for this paper we can find the following ones.
Definition 5.
Let $T$ be a t-norm and $I$ a fuzzy implication function. Then it is said that $I$ fulfills the
  • left neutrality principle:
    $I(1,y)=y,\quad\text{for all }y\in[0,1].$ ($)𝐍𝐏$
  • ordering property:
    $I(x,y)=1\Leftrightarrow x\leq y,\quad\text{for all }x,y\in[0,1].$ ($)𝐎𝐏$
  • $T$-conditionality with respect to $T$:
    $T(x,I(x,y))\leq y,\quad\text{for all }x,y\in[0,1].$ ($)𝐓𝐂$
Since Definition 3 is quite general, it allows the existence of many different families of fuzzy implication functions. Indeed, in [12] more than 100 families of fuzzy implication functions were gathered. Again, we only provide here the definition of those families related to the results of this paper.
Definition 6 ([15]).
A function $I:[0,1]^{2}\to[0,1]$ is called an $R$-implication if there exists a t-norm $T$ such that
$I(x,y)=\sup\{t\in[0,1]\mid T(x,t)\leq y\},\quad x,y\in[0,1].$
If $I$ is an $R$-implication generated from a t-norm $T$, then it is denoted by $I_{T}$.
Proposition 7 ([18, Definition 2.10]).
Let $C$ be a copula. The function $I_{C}:[0,1]^{2}\to[0,1]$ given by
$I_{C}(x,y)=\left\{\begin{array}[]{ll}1&\hbox{if }x=0,\\ \frac{C(x,y)}{x}&\hbox{if }x>0,\end{array}\right.$
is a fuzzy implication function if and only if
$C(x_{1},y)x_{2}\geq C(x_{2},y)x_{1},\quad\text{for all }x_{1}\leq x_{2}\text{ and }y\in[0,1].$
In this case, $I_{C}$ is called a probabilistic implication.
Definition 8 ([19, Definition 8]).
Let $k:[0,1]\to[0,1]$ be a strictly increasing and continuous function with $k(1)=1$. The function $I_{k}:[0,1]^{2}\to[0,1]$ defined by
$I_{k}(x,y)=k^{(-1)}\left(\frac{1}{x}\cdot k(y)\right),\quad x,y\in[0,1],$
with the understanding $\frac{0}{0}=1$ and $\frac{1}{0}=+\infty$ and where $k^{(-1)}$ is the pseudo-inverse of $k$, is called a $k$-generated implication. The function $k$ itself is called a $k$-generator.
III Fuzzy implicative rules and quality measures
In this section, we recall the modeling of fuzzy implicative rules proposed in [11] and we study which properties should the pair $(T,I)$ satisfy in order to obtain the desired behavior. Also, we provide the generalization of four of the most important quality measures (coverage, support, confidence, and weighted relative accuracy). Finally, we prove that our framework can be seen as the generalization of the crisp setting and fuzzy conjunctive rules proposals.
III-A Rule’s Modeling
We consider a set of $n_{f}\in\mathbb{N}$ features $\{X_{m}\mid m=1,\dots,n_{f}\}$, each with domain $D_{m}\subseteq\mathbb{R}$ and a target variable $Y$ with domain $D_{Y}\subseteq\mathbb{R}$. These variables can be categorical or numerical (including the target variable). A set of $n_{e}\in\mathbb{N}$ examples $E=\{E^{d}=(e_{1}^{d},\dots,e_{n_{f}}^{d},y^{d})\mid d=1,\dots,n_{e}\}$. For each feature $X_{m}\in\{X_{1},\dots,X_{n_{f}}\}$ we consider $l_{m}\in\mathbb{N}$ linguistic labels $X_{m}:\{LL_{m}^{1},\dots,LL_{m}^{l_{m}}\}$. For each feature $X_{m}\in\{X_{1},\dots,X_{n_{f}}\}$ and linguistic label $LL_{m}^{n}$, $n\in\{1,\dots,l_{m}\}$ we consider a membership function $\mu_{LL_{m}^{n}}:D_{m}\to[0,1]$. Moreover, for the target variable we consider $n_{c}\in\mathbb{N}$ linguistic labels that we call “classes” $Y:\{\text{Class}_{1},\dots,\text{Class}_{n_{c}}\}$. For each $\text{Class}_{j}$, $j\in\{1,\dots,n_{c}\}$ we consider a membership function $\mu_{Class_{j}}:D_{Y}\to[0,1]$.
For simplicity, we consider rules as the conjunction of literals, for other modelings an analogous study could be made. In this case, to determine a rule we fix a consequent $Class_{j}$ and a subset of features $S=\{X_{m_{1}},\dots,X_{m_{s}}\}\subseteq\{X_{1},\dots,X_{n_{f}}\}$ that will be considered in the antecedent and $L=\{LL^{n_{m_{1}}}_{m_{1}},\dots,LL^{n_{m_{s}}}_{m_{s}}\}$ the corresponding linguistic labels with $n_{m_{i}}\in\{1,\dots,l_{m}\}$ for all $X_{m}\in S$, i.e., only a linguistic label per considered feature in the rule. Then, a rule can be expressed as
$R^{S,L}_{j}:\text{IF }(X_{m_{1}}\text{ IS }LL_{m_{1}}^{n_{m_{1}}}\text{ AND }\dots\text{ AND }X_{m_{s}}\text{ IS }LL_{m_{s}}^{n_{m_{s}}})$
[…]
Now, to model the corresponding conjunction and conditional we consider a t-norm $T:[0,1]^{2}\to[0,1]$ and a fuzzy implication function $I:[0,1]^{2}\to[0,1]$, respectively. We will denote by $\mathcal{R}^{I,T}$ the set of all the rules where the conjunction is modeled by a t-norm $T$ and the conditional by a fuzzy implication function $I$, and by $\mathcal{S}(\mathcal{R}^{I,T})$ the set of all possible subsets of $\mathcal{R}^{I,T}$. Besides, we will denote by $\mathcal{R}^{I,T}_{j}$ the subset of $\mathcal{R}^{I,T}$ obtained by fixing Classj as the consequent and by $\mathcal{S}(\mathcal{R}_{j}^{I,T})$ the set of all possible subsets of $\mathcal{R}^{I,T}_{j}$.
Now, given a rule $R^{S,L}_{j}\in\mathcal{R}^{I,T}$ expressed as in Eq. (III-A) and an example $E^{d}\in E$ we define:
  • The truth value of the antecedent of the rule $R^{S,L}_{j}\in\mathcal{R}^{I,T}$ evaluated on example $E^{d}$ as
    $\mu_{ant}^{R^{S,L}_{j},E^{d}}=T\left(\mu_{LL_{m_{1}}^{n_{m_{1}}}}(e_{m_{1}}^{d}),\dots,\mu_{LL_{m_{s}}^{n_{m_{s}}}}(e_{m_{s}}^{d})\right).$
  • The truth value of the consequent of the rule $R^{S,L}_{j}\in\mathcal{R}^{I,T}$ evaluated on example $E^{d}$ as
    $\mu_{con}^{R^{S,L}_{j},E^{d}}=\mu_{Class_{j}}(y^{d}).$
  • The truth value of the rule $R^{S,L}_{j}\in\mathcal{R}^{I,T}$ evaluated on example $E^{d}$ as
    $\mu_{rule}^{R^{S,L}_{j},E^{d}}=I(\mu_{ant}^{R^{S,L}_{j},E^{d}},\mu_{con}^{R^{S,L}_{j},E^{d}}).$
  • The truth value of the evaluation of the rule $R^{S,L}_{j}\in\mathcal{R}^{I,T}$ evaluated on example $E^{d}$ as
    $\mu_{eval}^{R^{S,L}_{j},E^{d}}$ $=$ $T(\mu_{ant}^{R^{S,L}_{j},E^{d}},\mu_{rule}^{R^{S,L}_{j},E^{d}})$
    $=$ $T(\mu_{ant}^{R^{S,L}_{j},E^{d}},I(\mu_{ant}^{R^{S,L}_{j},E^{d}},\mu_{con}^{R^{S,L}_{j},E^{d}})).$
The properties of the fuzzy operators are of the utmost importance for the correct behavior of the rule mining technique [20]. If we do not pay attention to this issue we may encounter inconsistencies in the corresponding model. In accordance, we next discuss some desirable properties of fuzzy implicative rules and we compare them with the crisp case. First of all, let us introduce the concept of a refinement of a fuzzy rule as a new rule, whose antecedent has more restrictions.
Definition 9.
Let $I$ be a fuzzy implication function, $T$ a t-norm and $R^{\tilde{S},\tilde{L}}_{\tilde{j}}$, $R^{S,L}_{j}\in\mathcal{R}^{I,T}$. We say that $R^{\tilde{S},\tilde{L}}_{\tilde{j}}$ is a refinement of $R^{S,L}_{j}$ if and only if $\tilde{j}=j$, $S\subsetneq\tilde{S}\text{ and }LL_{m}^{n_{m}}\in\tilde{L}$ for all $X_{m}\in S$. In this case, we denote it by $R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{\tilde{j}}$.
In the crisp setting, the number of examples that fulfill the conditions in the antecedent is less in any refinement than in the original rule. In the fuzzy case this property is ensured by the monotonicity of the t-norm, and it is interpreted as follows: if $R^{\tilde{S},\tilde{L}}_{\tilde{j}}$ is a refinement of $R^{S,L}_{j}$, then for any example $E^{d}$ the truth value when evaluating it in the antecedent of $R^{\tilde{S},\tilde{L}}_{\tilde{j}}$ is smaller than in the antecedent of $R^{S,L}_{j}$.
Proposition 10.
Let $I$ be a fuzzy implication function, $T$ a t-norm and $Class_{j}\in\{Class_{1},\dots,Class_{n_{c}}\}$, then
$\mu_{ant}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}\leq\mu_{ant}^{R^{S,L}_{j},E^{d}},$
for all $R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{j},~{}R^{S,L}_{j},R^{\tilde{S},\tilde{L}}_{j}\in\mathcal{R}^{I,T}_{j}\text{ and }E^{d}\in E.$ Thus, $\sum_{d=1}^{n_{e}}\mu_{ant}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}\leq\sum_{d=1}^{n_{e}}\mu_{ant}^{R^{S,L}_{j},E^{d}}$.
Proof.
Without loss of generality let us consider $S=\{X_{m_{1}},\dots,X_{m_{s}}\}$, $L=\{LL_{m_{1}}^{n_{m_{1}}},\dots LL_{m_{s}}^{n_{m_{s}}}\}$ $\tilde{S}\setminus S=\{X_{m_{s+1}},\dots,X_{m_{r}}\}$ and $\tilde{L}\setminus L=\{LL_{m_{s+1}}^{n_{m_{s+1}}},\dots LL_{m_{r}}^{n_{m_{r}}}\}$. By the monotonicity of the t-norm $T$ we have
$\mu_{ant}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}$ $=$ $T\left(\mu_{LL_{m_{1}}^{n_{m_{1}}}}(e_{m_{1}}^{d}),\dots,\mu_{LL_{m_{s}}^{n_{m_{s}}}}(e_{m_{s}}^{d}),\right.$
$\left.\mu_{LL_{m_{s+1}}^{n_{m_{s+1}}}}(e_{m_{s+1}}^{d}),\dots,\mu_{LL_{m_{r}}^{n_{m_{r}}}}(e_{m_{r}}^{d})\right)$
$\leq$ $T\left(\mu_{LL_{m_{1}}^{n_{m_{1}}}}(e_{m_{1}}^{d}),\dots,\mu_{LL_{m_{s}}^{n_{m_{s}}}}(e_{m_{s}}^{d})\right)$
$=$ $\mu_{ant}^{R^{S,L}_{j},E^{d}}.\hskip 113.81102pt\qed$
On the other hand, in a crisp rule the number of examples that fulfill the conditions in the antecedent of a rule and also belong to the target class is smaller than number of the examples that belong to the target class. In the case of fuzzy rules the analogous fact is not straightforward. Since we have generalized the concept of fulfilling the antecedent and belonging to the target class in terms of the generalized modus ponens, we have to impose the $T$-conditionality to ensure that the truth value when evaluating an example in a rule is smaller than the membership degree of the example to the consequent.
Proposition 11.
Let $I$ be a fuzzy implication function and $T$ a t-norm. If $I$ satisfies (TC) with respect to $T$ then
$\mu_{eval}^{R^{S,L}_{j},E^{d}}\leq\mu_{con}^{R^{S,L}_{j},E^{d}},\quad\text{forall }R^{S,L}_{j}\in\mathcal{R}^{I,T}\text{ and }E^{d}\in E.$
Thus, $\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{S,L}_{j},E^{d}}\leq\sum_{d=1}^{n_{e}}\mu_{con}^{R^{S,L}_{j},E^{d}}$.
Proof.
By (TC) we have
$\mu_{eval}^{R^{S,L}_{j},E^{d}}=T(\mu_{ant}^{R^{S,L}_{j},E^{d}},I(\mu_{ant}^{R^{S,L}_{j},E^{d}},\mu_{con}^{R^{S,L}_{j},E^{d}}))\leq\mu_{con}^{R^{S,L}_{j},E^{d}}.$
Besides, in crisp rules the number of examples that fulfill the conditions in the antecedent of a rule and also belong to the target class is smaller in any refinement than in the original rule. This is because in any refinement there are more conditions in the antecedent, so it is more complex. Again, in the case of fuzzy rules we do not generally have the analogous property. However, differently from the previous case, in order to obtain the desired property for fuzzy rules we have to impose an additional property of fuzzy implication functions for which we could not find any study about it in the consulted bibliography. This property captures the intuitive idea that the truth value of the inference obtained by applying the generalized modus ponens should be decreasing with respect to the truth value of the antecedent.
Definition 12.
Let $I$ be a fuzzy implication function and $T$ a t-norm. We say that $I$ satisfies the monotonicity of the generalized modus ponens with respect to $T$ if and only if
$T(\tilde{x},I(\tilde{x},y))\leq T(x,I(x,y)),\quad x,\tilde{x},y\in[0,1]\text{ with }\tilde{x}\leq x.$ (MTC)
Now, if we impose (MTC) we can ensure that the truth value of the evaluation of an example is smaller in any refinement than in the original rule.
Proposition 13.
Let $I$ be a fuzzy implication function, $T$ a t-norm and $Class_{j}\in\{Class_{1},\dots,Class_{n_{c}}\}$. If $I$ satisfies (MTC) with respect to $T$ then
$\mu_{eval}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}\leq\mu_{eval}^{R^{S,L}_{j},E^{d}},$
for $\text{for all }R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{j},~{}R^{S,L}_{j},R^{\tilde{S},\tilde{L}}_{j}\in\mathcal{R}^{I,T}_{j}\text{ and }E^{d}\in E.$ Thus, $\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}\leq\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{S,L}_{j},E^{d}}$.
Proof.
Let us consider $R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{j}$, by Proposition 10 we know that $\mu_{ant}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}\leq\mu_{ant}^{R^{S,L}_{j},E^{d}}$, and since the two rules consider the same target class we have $\mu_{con}^{R^{\tilde{S},\tilde{L}}_{j}}=\mu_{con}^{R^{S,L}_{j},E^{d}}$. Thus, by (MTC) we obtain
$\mu_{eval}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}$ $=$ $T(\mu_{ant}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}},I(\mu_{ant}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}},\mu_{con}^{R^{S,L}_{j},E^{d}}))$
$\leq$ $T(\mu_{ant}^{R^{S,L}_{j},E^{d}},I(\mu_{ant}^{R^{S,L}_{j},E^{d}},\mu_{con}^{R^{S,L}_{j},E^{d}}))$
$=$ $\mu_{eval}^{R^{S,L}_{j},E^{d}}.$
Notice that the property (MTC) was not considered in [11], so in that approach it was not guaranteed that the evaluation of a rule was monotone with respect to our definition of refinement of a rule (see Definition 9).
Finally, in view of the discussion above, we may conclude that the pair of operators $(T,I)$ should satisfy (TC) and (MTC) in order to behave adequately when used for fuzzy implicative rule mining.
Definition 14.
Let $I$ be a fuzzy implication function and $T$ a t-norm. We say that the pair $(T,I)$ is adequate for fuzzy implicative rule mining if $I$ fulfills (TC) and (MTC) with respect to $T$.
In the subsequent sections, we prove that these two properties are not only necessary for the reasons pointed out above, but are also crucial for ensuring other desired properties.
III-B Fuzzy implicative rules as a generalization of others
Let us point out that the interpretation of the logical conditional relies completely on the selection of the corresponding fuzzy implication function. Thus, our framework is very flexible. Indeed, in this section we prove that the rule’s modeling proposed in Section III-A can be seen as the generalization of the crisp setting and other fuzzy logic perspectives based on conjunctive fuzzy rules.
In the crisp setting the target variable is categorical and, in this case, we can construct a bijection between the domain of the target variable $D_{Y}$ with $|D_{Y}|=n_{c}$ and the set $\{1,\dots,n_{c}\}$, so let us assume $D_{Y}=\{1,\dots,n_{c}\}$. In this situation, it is clear that we are forced to construct a fuzzy set with a singleton membership function for each possible value of $D_{Y}$, so for all $j\in D_{Y}$ we consider $Class_{j}=j$ and
$\begin{array}[]{rcl}\mu_{Class_{j}}:D_{Y}&\longrightarrow&\{0,1\}\\ \tilde{j}&\longmapsto&\left\{\begin{array}[]{ll}1&\text{if }\tilde{j}=j,\\ 0&\text{if }\tilde{j}\not=j.\end{array}\right.\end{array}$
Then, it is obvious that when evaluating a rule for a certain example the only possible values for the consequent are 00 or $1$, so the only values of the fuzzy implication function that are being used are $I(x,1)$ and $I(x,0)$ for all $x\in[0,1]$. Since $I(x,1)=1$ for all $x\in[0,1]$, only the choice of the natural negation $N_{I}$ plays a role in this framework. However, if we take into account that in the end we are interested in the evaluation of the rule, if we choose a fuzzy implication function $I$ which satisfies (TC) with respect to $T$ we have that
$T(x,I(x,0))\leq 0\Rightarrow T(x,I(x,0))=0.$
Thus, when the consequent is zero the evaluation of the rule is also zero independently from the natural negation of the corresponding fuzzy implication function. In accordance, when a singleton membership function for the target variable is considered, the role of the fuzzy implication function disappears and the truth value of the antecedent of the rule is considered as the truth value of the evaluation of the rule, whenever the consequent is non-zero. Then, by imposing the restriction ${\bf(TC)}$ we generalize other perspectives of fuzzy rules where neither numeric targets nor fuzzy implication functions were considered (see for instance [21]).
Proposition 15.
Let $I$ be a fuzzy implication function, $T$ a t-norm and $Class_{j}\in\{Class_{1},\dots,Class_{n_{c}}\}$ such that $\mu_{Class_{j}}:D_{Y}\to\{0,1\}$. If $I$ satisfies (TC) with respect to $T$ then $\mu_{con}^{R^{S,L}_{j},E^{d}}\in\{0,1\}$ and
$\mu_{rule}^{R^{S,L}_{j},E^{d}}=\left\{\begin{array}[]{ll}N_{I}(\mu_{ant}^{R^{S,L}_{j},E^{d}})&\text{if }\mu_{con}^{R^{S,L}_{j},E^{d}}=0,\\[10.00002pt] 1&\text{if }\mu_{con}^{R^{S,L}_{j},E^{d}}=1,\end{array}\right.$
$\mu_{eval}^{R^{S,L}_{j},E^{d}}=\left\{\begin{array}[]{ll}0&\text{if }\mu_{con}^{R^{S,L}_{j},E^{d}}=0,\\[10.00002pt] \mu_{ant}^{R^{S,L}_{j},E^{d}}&\text{if }\mu_{con}^{R^{S,L}_{j},E^{d}}=1,\end{array}\right.$
for all $R^{S,L}_{j}\in\mathcal{R}^{I,T}_{j}\text{ and }E^{d}\in E$.
Proof.
Straightforward. ∎
On the other hand, there are other fuzzy logic perspectives that consider fuzzy linguistic variables also for the consequent of the rules but the evaluation of the rules is performed by using a t-norm [20], i.e., $\mu_{eval}^{R^{S,L}_{j},E^{d}}=T(\mu_{ant}^{R^{S,L}_{j},E^{d}},\mu_{con}^{R^{S,L}_{j},E^{d}})$. In this case, it is clear that if we select the following fuzzy implication function
$I_{Y}(x,y)=\left\{\begin{array}[]{ll}1&\text{if }x=0\text{ or }y=1,\\[10.00002pt] y&\text{otherwise},\end{array}\right.$ (1)
then our approach is equivalent.
Proposition 16.
Let $T$ be a t-norm and $I_{Y}$ the fuzzy implication function in Eq. (1). Then, $I_{Y}$ satisfies (TC) and (MTC) with respect to $T$ and
$\mu_{eval}^{R^{S,L}_{j},E^{d}}=T(\mu_{ant}^{R^{S,L}_{j},E^{d}},\mu_{con}^{R^{S,L}_{j},E^{d}}),$
for all $R^{S,L}_{j}\in\mathcal{R}^{I,T}_{j}\text{ and }E^{d}\in E$.
Proof.
Let $T$ be a t-norm and $I_{Y}$ the fuzzy implication function Eq. (1), then
$T(x,I_{Y}(x,y))$ $=$ $\left\{\begin{array}[]{ll}x&\text{if }x=0\text{ or }y=1,\\[10.00002pt] T(x,y)&\text{otherwise},\end{array}\right.$
$=$ $T(x,y).$
III-C Quality measures
One of the key decisions when designing a rule mining algorithm is how to select and order the rules that are more interesting for a concrete goal. Since the goal may not be the same depending on the task or the desires of an expert, a plethora of quality measures have been considered (see [22] for an overview). In fact, a quality measure can be generally defined as any function $q:\mathcal{R}^{I,T}\to\mathbb{R}$ which assigns a numeric value to each rule $R^{S,L}_{j}\in\mathcal{R}^{I,T}$ [9].
In this section, we give a generalization of four of the most widely studied quality measures in rule mining algorithms: Coverage, Support, Confidence and Unusualness or Weighted Relative Accuracy (WRAcc). In our approach, the overall truth values of the antecedent, the consequent and the evaluation of the rule are used for the computation of the quality measures instead of the corresponding percentages of examples. Although some of these quality measures were already proposed in [11] and also in other fuzzy perspectives that do not take into account fuzzy implication functions [23, 20], in this section we have adapted their definition to our framework and also we have provided a generalization of the Weighted relative accuracy measure. Further, we justify the correct behavior of these measures with respect to the selected fuzzy operators.
Definition 17.
Let $I$ be a fuzzy implication function and $T$ a t-norm, we define the fuzzy coverage as the function $FCov:\mathcal{R}^{I,T}\to[0,1]$ given by
$FCov(R^{S,L}_{j})=\frac{\displaystyle\sum_{d=1}^{n_{e}}\mu_{ant}^{R^{S,L}_{j},E^{d}}}{|E|},$
and it measures the average truth value of the antecedent for all the examples.
Definition 18.
Let $I$ be a fuzzy implication function and $T$ a t-norm, we define the fuzzy support as the function $FSupp:\mathcal{R}^{I,T}\to[0,1]$ given by
$FSupp(R^{S,L}_{j})=\frac{\displaystyle\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{S,L}_{j},E^{d}}}{|E|},$
and it measures the average truth value of the evaluation of the rule for all the examples.
Definition 19.
Let $I$ be a fuzzy implication function and $T$ a t-norm, we define the fuzzy confidence as the function $FConf:\mathcal{R}^{I,T}\to[0,1]$ given by
$FConf(R^{S,L}_{j})=\frac{FSupp(R^{S,L}_{j})}{FCov(R^{S,L}_{j})}=\frac{\displaystyle\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{S,L}_{j},E^{d}}}{\displaystyle\sum_{d=1}^{n_{e}}\mu_{ant}^{R^{S,L}_{j},E^{d}}},$
and it measures the quotient of the overall truth value of the evaluation of the rule and the overall truth value of the antecedent for all examples.
Definition 20.
Let $I$ be a fuzzy implication function and $T$ a t-norm, we define the fuzzy weighted relative accuracy as the function $FWRAcc:\mathcal{R}^{I,T}\to[-1,1]$ given by
$FWRAcc(R^{S,L}_{j})=$
$FCov(R^{S,L}_{j})\cdot\left(FConf(R^{S,L}_{j})-\frac{\displaystyle\sum_{d=1}^{n_{e}}\mu_{con}^{R^{S,L}_{j},E^{d}}}{|E|}\right),$
and it measures the balance between the fuzzy coverage of the rule and its accuracy gain which is computed as the difference between the fuzzy confidence and the average truth value of the consequent for all the examples.
It is clear that the above quality measures have different goals [22]: the support and coverage are measures of generality since they quantify rules according to the individual patterns of interest covered; the confidence is a measure of precision; the unusualness is a measure of interest since it is designed to quantify the potential interest for the user.
Now, let us point out that the measure of fuzzy coverage is monotone with respect to the refinements, i.e., any refinement of a rule has a lower fuzzy coverage. This fact is also true for the fuzzy support if we impose that the pair ($T$,$I$) satisfies (MTC). This highlights again the importance of the newly introduced property. If we do not consider (MTC), the fuzzy support will not be necessarily monotone, and then it will not be adequate as a measure of generality.
Proposition 21.
Let $I$ be a fuzzy implication function, $T$ a t-norm and $Class_{j}\in\{Class_{1},\dots,Class_{n_{c}}\}$, then
$FCov(R^{S,L}_{j})\geq FCov(R^{\tilde{S},\tilde{L}}_{j}),$
for all $R^{S,L}_{j},R^{\tilde{S},\tilde{L}}_{j}\in\mathcal{R}^{I,T}_{j}$ such that $R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{j}$.
Proof.
Straightforward from Proposition 10. ∎
Proposition 22.
Let $I$ be a fuzzy implication function, $T$ a t-norm and $Class_{j}\in\{Class_{1},\dots,Class_{n_{c}}\}$. If $I$ satisfies (MTC) with respect to $T$ then
$FSupp(R^{S,L}_{j})\geq FSupp(R^{\tilde{S},\tilde{L}}_{j}),$
for all $R^{S,L}_{j},R^{\tilde{S},\tilde{L}}_{j}\in\mathcal{R}^{I,T}$ such that $R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{j}$.
Proof.
Straightforward from Proposition 13. ∎
IV Solutions for adequate fuzzy operators
In Section III we point out that a pair $(T,I)$ of t-norm and fuzzy implication function used for modeling fuzzy implicative rules should satisfy two properties: (TC) and (MTC). It is clear that there exist families of fuzzy implication functions that satisfy ${\bf(TC)}$, but since we have not found any information about (MTC) in the literature, in order to select an adequate pair $(T,I)$ we have to further study this property.
It is well known that the main families of fuzzy implication functions satisfy (NP) [17]. In accordance, we first prove that if $I$ satisfies (NP) then (MTC) implies (TC). Thus, for families that satisfy (NP) we only need to study the new property (MTC) for the available solutions of (TC).
Proposition 23.
Let $I$ be a fuzzy implication function and $T$ a t-norm. If $I$ satisfies $I(1,y)\leq y$ for all $y\in[0,1]$ and (MTC) with respect to $T$, then $I$ also satisfies (TC) with respect to $T$. In particular, (NP) and (MTC) with respect to $T$ imply (TC) with respect to $T$.
Proof.
Let $x\in[0,1]$, then by (MTC)
$T(x,I(x,y))\leq T(1,I(1,y))=I(1,y)\leq y.$
Next, we prove that if a fuzzy implication function satisfies (OP) then the verification of (MTC) can be simplified to the study of the monotonicity of a family of unary functions.
Proposition 24.
Let $I$ be a fuzzy implication function and $T$ a t-norm. If $I$ satisfies (OP) then $I$ satisfies (MTC) if and only if the function
$\begin{array}[]{rcl}f_{y}:[y,1]&\longrightarrow&[0,1]\\ x&\longmapsto&T(x,I(x,y))\end{array}$
is increasing for all $y\in[0,1)$.
Proof.
Let $I$ be a fuzzy implication function that satisfies (OP) and $T$ a t-norm such that $f_{y}$ is an increasing function for all $y\in[0,1)$. Let us prove that the pair $(T,I)$ satisfies (MTC) by distinguishing between three cases.
  • If $y\geq x\geq\tilde{x}$ then by (OP)
    $T(\tilde{x},I(\tilde{x},y))$ $=$ $T(\tilde{x},1)=\tilde{x}\leq x=T(x,1)$
    $=$ $T(x,I(x,y)).$
  • If $x\geq\tilde{x}\geq y$ then since $f_{y}$ is increasing
    $T(\tilde{x},I(\tilde{x},y))=f_{y}(\tilde{x})\leq f_{y}(x)\leq T(x,I(x,y)).$
  • If $x\geq y\geq\tilde{x}$ then by (OP) and since $f_{y}$ is increasing
    $T(\tilde{x},I(\tilde{x},\tilde{y}))$ $=$ $T(\tilde{x},1)=\tilde{x}\leq y=T(y,1)$
    $=$ $T(y,I(y,y))=f_{y}(y)\leq f_{y}(x)$
    $=$ $T(x,I(x,y)).$
If $I$ satisfies (MTC) it is straightforward to verify that the function $f_{y}$ is increasing for all $y\in[0,1)$. ∎
Having said this, we first gather some results about four families in the literature that satisfy (TC). The families considered in this section have been: the $R$-implications with a left-continuous t-norms [17, Section 2.5] as one of the most-well known families of fuzzy implication functions (which were originally considered in [11] for modelling fuzzy implicative association rules); the probabilistic implications which interpret the probability of an implication as the conditional probability [18]; $k$-implications which are generated by multiplicative generators of continuous Archimedean t-norms [19]; and strict/nilpotent $T$-power invariant implications as a family of fuzzy implication functions which satisfy the invariance with respect to the powers of a certain t-norm [24, 25]. The selection of these four families among the great variety available has been twofold: we needed a family that has constructive solutions of the $T$-conditionality for which is relatively easy to later study (MTC) and also that has a nice structure for applications. In future studies, the families considered could be enlarged.
Corollary 25.
The following statements hold:
  1. (i)
    Let $T$ be a left-continuous Archimedean t-norm and $I_{T}$ its $R$-implication. Then $I_{T}$ satisfies (TC) with $T$.
  2. (ii)
    Let $I_{C}$ be a probabilistic implication, then $I_{C}$ satisfies (TC) with $T_{\bm{P}}$.
  3. (iii)
    Let $k$ be a $k$-generator with $k\leq\text{id}_{[0,1]}$, $I_{k}$ the $k$-generated implication and $T_{k}$ the t-norm generated by $k$ as its multiplicative generator. Then, $I_{k}$ satisfies (TC) with each t-norm $T$ that is weaker than $T_{k}$, i.e., $T\leq T_{k}$.
  4. (iv)
    Let $I^{T}_{\varphi,f,g}$ be a strict $T$-power invariant implication. Then $I^{T}_{\varphi,f,g}$ satisfies (TC) with respect to $T$ if and only if $\varphi(w)=0$ for all $w\in[0,1)$, and $f(x)=g(y)=0$ for all $x,y\in(0,1)$. In this case, $I^{T}_{\varphi,f,g}$ satisfies (TC) with respect to any t-norm $T^{*}$.
  5. (v)
    Let $I^{T}_{\varphi,f,g}$ be a nilpotent $T$-power invariant implication and $t$ an additive generator of $T$. Then $I^{T}_{\varphi,f,g}$ satisfies (TC) with respect to $T$ if and only if $\varphi(w)\leq t^{-1}\left(t(0)(1-w)\right)$ for all $w\in[0,1)$, $f(x)\leq t^{-1}(t(0)-t(x))$ for all $x\in(0,1)$ and $g(y)=0$ for all $y\in(0,1)$.
Proof.
  1. (i)
    [17, Theorem 7.4.8].
  2. (ii)
    [18, Proposition 4.3].
  3. (iii)
    [19, Proposition 11].
  4. (iv)
    [26, Proposition 4.44].
  5. (v)
    [26, Proposition 4.73].
Next, we point out particular cases of fuzzy implication functions that belong to one of these five families and satisfy (MTC). Also, we consider the fuzzy implication function in Eq. (1) just to recall again that this operator generalizes the fuzzy conjunctive rule perspective.
Proposition 26.
The following statements hold:
  1. (i)
    The pair $(I_{\bm{GD}},T_{\bm{M}})$ satisfies (MTC). In this case,
    $T_{\bm{M}}(x,I_{\bm{GD}}(x,y))=\min\{x,y\}.$
  2. (ii)
    Let $T$ be a left-continuous Archimedean t-norm and $I_{T}$ its $R$-implication. Then $I_{T}$ satisfies (MTC) with $T$. In this case,
    $T(x,I_{T}(x,y))=\min\{x,y\}.$
  3. (iii)
    Let $I_{C}$ be a probabilistic implication, then $I_{C}$ satisfies (MTC) with $T_{\bm{P}}$. In this case,
    $T_{\bm{P}}(x,I_{C}(x,y))=C(x,y).$
  4. (iv)
    Let $k$ be a $k$-generator, $I_{k}$ the $k$-generated implication and $T_{k}$ the t-norm generated by $k$ as its multiplicative generator. If $k(\tilde{x})\cdot x\leq k(x)\cdot\tilde{x}$ for all $0\leq\tilde{x}\leq x\leq 1$, then $I_{k}$ satisfies (MTC) with $T_{k}$. In this case,
    $T_{k}(x,I_{k}(x,y))=k^{-1}\left(k(x)\cdot\min\left\{\frac{k(y)}{x},1\right\}\right).$
  5. (v)
    Let $T$ be a continuous Archimedean t-norm and $I^{T}_{\varphi,f,g}$ the strict/nilpotent $T$-power invariant implication obtained from it. If $I^{T}_{\varphi,f,g}$ satisfies (TC) and (MTC) with $T$ then $T(x,I^{T}_{\varphi,f,g}(x,y))=0$ for all $x\in[0,1]$ and $y\in(0,1)$.
  6. (vi)
    Let $T$ be any t-norm and $I_{Y}$ the fuzzy implication function given by Eq. (1), then $I_{Y}$ satisfies (MTC) with respect to $T$. In this case,
    $T(x,I_{Y}(x,y))=T(x,y).$
Proof.
  1. (i)
    Let $y\in[0,1)$, then
    $T_{\bm{M}}(x,I_{\bm{GD}}(x,y))$ $=$ $\left\{\begin{array}[]{ll}T_{\bm{M}}(x,1)&\text{if }x\leq y,\\ T_{\bm{M}}(x,y)&\text{if }x>y,\end{array}\right.$
    $=$ $\left\{\begin{array}[]{ll}x&\text{if }x\leq y,\\ y&\text{if }x>y,\end{array}\right.$
    and since $I_{\bm{GD}}$ satisfies (OP), by Proposition 24 we have that $I_{\bm{GD}}$ satisfies (MTC) with $T_{\bm{M}}$.
  2. (ii)
    Let $T$ be a continuous Archimedean t-norm and $t$ an additive generator of $T$, then the corresponding $R$-implication is given by
    $I_{T}(x,y)=\left\{\begin{array}[]{ll}1&\text{if }x\leq y,\\ t^{-1}(t(y)-t(x))&\text{if }x>y.\end{array}\right.$
    Let $y\in[0,1)$, then
    $T(x,I_{T}(x,y))$ $=$ $\left\{\begin{array}[]{ll}T(x,1)&\text{if }x\leq y,\\ T(x,t^{-1}(t(y)-t(x)))&\text{if }x>y,\end{array}\right.$
    $=$ $\left\{\begin{array}[]{ll}x&\text{if }x\leq y,\\ t^{(-1)}(t(x)+t(y)-t(x))&\text{if }x>y,\end{array}\right.$
    $=$ $\left\{\begin{array}[]{ll}x&\text{if }x\leq y,\\ y&\text{if }x>y.\end{array}\right.$
    Since $I_{T}$ satisfies (OP), by Proposition 24 we have that $I_{T}$ satisfies (MTC) with $T$.
  3. (iii)
    Let $I_{C}$ be a probabilistic implication, then for all $\tilde{x},x,y\in[0,1]$ such that $0<\tilde{x}\leq x$ we have
    $T_{\bm{P}}(\tilde{x},I_{C}(\tilde{x},y))$ $=$ $\tilde{x}\cdot\frac{C(\tilde{x},y)}{\tilde{x}}=C(\tilde{x},y)\leq C(x,y)$
    $=$ $T_{\bm{P}}(x,I_{C}(x,y)).$
    On the other hand, if $\tilde{x}=0$ then $T_{\bm{P}}(\tilde{x},I_{C}(\tilde{x},y))=T_{\bm{P}}(0,I_{C}(0,y))=0\leq T_{\bm{P}}(x,I_{C}(x,y))$ for all $x\in[0,1]$.
  4. (iv)
    According to the proof of [19, Proposition 11] we have
    $T_{k}(x,I_{k}(x,y))=k^{-1}\left(k(x)\cdot\min\left\{\frac{k(y)}{x},1\right\}\right).$
    Let $\tilde{x},x,y\in[0,1]$, if $\tilde{x}=0$ then it is clear that $T_{k}(\tilde{x},I_{k}(\tilde{x},y))=T_{k}(0,I_{k}(0,y))=0\leq T_{k}(x,I_{k}(x,y))$, so let us consider $0<\tilde{x}\leq x\leq 1$. We distinguish between different cases:
    • If $k(y)>x$ then $k(y)>x\geq\tilde{x}$ and
      $T_{k}(\tilde{x},I_{k}(\tilde{x},y))=\tilde{x}\leq x=T_{k}(x,I_{k}(x,y)).$
    • If $k(y)>\tilde{x}$ and $k(y)\leq x$, since $\frac{k(\tilde{x})}{\tilde{x}}\leq\frac{k(x)}{x}$ we have
      $T_{k}(x,I_{k}(x,y))$ $=$ $k^{-1}\left(\frac{k(x)k(y)}{x}\right)$
      $\geq$ $k^{-1}\left(\frac{k(\tilde{x})k(y)}{\tilde{x}}\right)\geq k^{-1}(k(\tilde{x}))$
      $=$ $\tilde{x}=T_{k}(\tilde{x},I_{k}(\tilde{x},y)).$
    • If $k(y)\leq\tilde{x}$, since $\frac{k(\tilde{x})}{\tilde{x}}\leq\frac{k(x)}{x}$ we have
      $T_{k}(\tilde{x},I_{k}(\tilde{x},y))$ $=$ $k^{-1}\left(\frac{k(\tilde{x})k(y)}{\tilde{x}}\right)$
      $\leq$ $k^{-1}\left(\frac{k(x)k(y)}{x}\right)$
      $=$ $T_{k}(x,I_{k}(x,y)).$
  5. (v)
    Let $T$ be a continuous Archimedean t-norm and $I^{T}_{\varphi,f,g}$ the corresponding strict/nilpotent $T$-power invariant implication. If $I^{T}_{\varphi,f,g}$ satisfies (TC) with respect to $T$ then by [26, Propositions 4.44 and 4.73] we obtain that $\varphi(0^{+})=0$ and by [24, Proposition 3] and [25, Proposition 20] we have that $I^{T}_{\varphi,f,g}(1,y)=g(y)=0$ for all $y\in(0,1)$. Then, since $I^{T}_{\varphi,f,g}$ satisfies (MTC) with $T$ we have
    $T(x,I^{T}_{\varphi,f,g}(x,y))$$\leq$ $T(1,I^{T}_{\varphi,f,g}(1,y))=T(1,g(y))$
    $=$ $T(1,0)=0,$
    which implies $T(x,I^{T}_{\varphi,f,g}(x,y))=0$ for all $x\in[0,1]$ and $y\in(0,1)$.
  6. (vi)
    It is proved in Proposition 16.
From Proposition 26 we can conclude the following:
  • If $T$ is the minimum or a continuous Archimedean t-norm then it satisfies (MTC) with the corresponding $R$-implication, $I_{T}$. Moreover, by Corollary 25 we know that in this case $(T,I_{T})$ also satisfies (TC). Thus, we deduce that $(T,I_{T})$ is an adequate pair for fuzzy implicative rule mining when $T$ is the minimum or a continuous Archimedean t-norm. Notice that we have not studied (MTC) when $T$ is different from the minimum or a continuous Archimedean t-norm. These cases could be performed in future studies of this property.
  • If $I_{C}$ is a probabilistic implication, then it satisfies (MTC) with respect to the product t-norm. Thus, $(T_{\bm{P}},I_{C})$ is an adequate pair for fuzzy implicative rule mining.
  • If $k$ is a $k$-generator with $k(\tilde{x})\cdot x\leq k(x)\cdot\tilde{x}$ for all $0\leq\tilde{x}\leq x\leq 1$ then $(T_{k},I_{k})$ where $I_{k}$ is the $k$-generated implication and $T_{k}$ the t-norm generated by $k$ as its multiplicative generator satisfy (MTC). In this case, if we consider $x=1$ then we have $k(\tilde{x})\leq k(1)\cdot\tilde{x}=\tilde{x}=\text{id}(\tilde{x})$ for all $0\leq\tilde{x}\leq 1$ and the pair $(T_{k},I_{k})$ also satisfies (TC). In particular, $k(0)=0$ in this case. Then, $(T_{k},I_{k})$ is an adequate pair for fuzzy implicative rule mining under these conditions. For instance, let us consider the following $k$-generators:
    $k_{\lambda}^{SS}(x)=e^{\frac{x^{\lambda}-1}{\lambda}},\quad\lambda\in(-\infty,0),$
    $k_{\lambda}^{H}(x)=\frac{x}{\lambda+(1-\lambda)x},\quad\lambda\in(1,+\infty),$
    $k_{\lambda}^{F}(x)=\frac{\lambda^{x}-1}{\lambda-1},\quad\lambda\in(1,+\infty).$
    Then, we consider the functions $f_{k_{\lambda}}^{L}(x)=\frac{k_{\lambda}^{L}(x)}{x}$ for all $x\in(0,1]$ and $L\in\{SS,H,F\}$ and we compute their first derivative:
    $(f_{k_{\lambda}}^{SS})^{\prime}(x)=\frac{e^{\frac{x^{\lambda}-1}{\lambda}}\cdot(x^{\lambda}-1)}{x^{2}},\quad\lambda\in(-\infty,0),$
    $(f_{k_{\lambda}}^{H})^{\prime}(x)=\frac{\lambda-1}{(\lambda+(1-\lambda)x)^{2}},\quad\lambda\in(1,+\infty),$
    $(f_{k_{\lambda}}^{F})^{\prime}(x)=\frac{\lambda^{x}(x\ln\lambda-1)+1}{x^{2}(\lambda-1)},\quad\lambda\in(1,+\infty).$
    It is straightforward to verify that $(f_{k_{\lambda}}^{L})^{\prime}(x)>0$ for all $x\in(0,1]$, so all $f_{k_{\lambda}}^{L}$ are increasing and consequently, $k_{\lambda}^{L}(\tilde{x})\cdot x\leq k_{\lambda}^{L}(x)\cdot\tilde{x}$ for all $0\leq\tilde{x}\leq x\leq 1$, $L\in\{SS,H,F\}$ and the corresponding domain of $\lambda$. Therefore, all the pairs $(T_{k_{\lambda}^{L}},I_{k_{\lambda}^{L}})$ are adequate for fuzzy implicative rule mining (in [19, Example 1] the reader can find the corresponding constructions).
  • The point (v)-Proposition 26 discloses that the family of $T$-power invariant implications does not provide interesting solutions when studying the property (MTC). Indeed, any solution of this property returns zero when applying the generalized modus ponens. Thus, for the moment there are no solutions if we want to use fuzzy implication functions which are invariant with respect to hedges modeled through powers of t-norms.
  • Finally, if we consider any t-norm and the fuzzy implication function in Eq. (1) then we obtain a pair of operators which is adequate for fuzzy implicative rule mining and the modeling is equivalent to fuzzy conjunctive rules.
Summarizing the above discussion, in the following proposition we point out five different pairs of fuzzy implication function and t-norm that are adequate for fuzzy implicative rule mining. More specifically, in Table I we propose four concrete examples of adequate pairs which are the ones considered in our experimental results.
Proposition 27.
The following pairs are adequate for fuzzy implicative rule mining in the sense of Definition 14:
  • $(T_{\bm{M}},I_{\bm{GD}})$.
  • $(T,I_{\bf T})$ where $T$ is a continuous, Archimedean t-norm.
  • $(T_{\bm{P}},I_{C})$ where $I_{C}$ is a probabilistic fuzzy implication.
  • $(T_{k},I_{k})$ where $k$ is a $k$-generator with $k(\tilde{x})\cdot x\leq k(x)\cdot\tilde{x}$ for all $0\leq\tilde{x}\leq x\leq 1$, $I_{k}$ is a $k$-generated implication and $T_{k}$ the t-norm generated by $k$ as its multiplicative generator.
  • $(T,I_{Y})$ where $I_{Y}$ is the fuzzy implication function in Eq. (1) and $T$ is any t-norm.
Once we have several options for the selection of the pair of operators $(T,I)$ it is interesting to look at the structure of the evaluation of the antecedent and the rule for different pairs. In Proposition 26 we can see the expression of the truth value of the evaluation of the rule when the generalized modus ponens is used for each case. Notice that even though fuzzy implication functions are non-commutative, it may happen that the expression given by the generalized modus ponens is commutative. Indeed, this is the case for $R$-implications and also probabilistic implications that use a commutative copula. The user might like to avoid these options if the purpose is to obtain directional rules, in which it is significantly different the permutation of antecedent and consequent. That is why in the last two rows in Table I we have considered pairs of operators that result in non-commutative evaluations of the corresponding rules by considering a non-commutative copula for a probabilistic implication and a $k$-implication.
TABLE I: Four examples of pairs $(T,I)$ which fulfill (TC) and (MTC).
t-norm $\bm{T}$ Fuzzy implication function $\bm{I}$
Product t-norm
$T_{\bm{P}}(x,y)=xy$.
$I_{Y}(x,y)=\left\{\begin{array}[]{ll}1&\text{if }x=0\text{ and }y=1,\\ y&\text{otherwise}.\end{array}\right.$
Łukasiewicz t-norm
$T_{\bm{LK}}(x,y)=\max\{x+y-1,0\}$.
Łukasiewicz fuzzy implication
$I_{\bm{LK}}(x,y)=\min\{1,1-x+y\}.$
Product t-norm
$T_{\bm{P}}(x,y)=xy$.
Probabilistic fuzzy implication based on
$(C_{\lambda}^{H})_{1,0}$ copula
$I_{\bm{(C_{\lambda}^{H})_{1,0}}}(x,y)=$
$\par\left\{\begin{array}[]{ll}1&\text{ if }x=0,\\ \frac{y}{x}-\frac{(1-x)y}{x(\lambda+(1-\lambda)(1-x+xy))}&\text{ if }x>0,\end{array}\right.$
with $\lambda\in[0,2]$. Considered value $\lambda=0.01$.
Schweizer-Sklar t-norm
$T_{\lambda}^{\bm{SS}}(x,y)=(\max\{(x^{\lambda}+y^{\lambda}-1),0\})^{\frac{1}{\lambda}}.$
$k_{\lambda}$-generated Schweizer-Sklar implications
$I^{\bm{SS}}_{k_{\lambda}}(x,y)=\left\{\begin{array}[]{ll}1&\text{ if }x\leq e^{\frac{y^{\lambda}-1}{\lambda}},\\ (y^{\lambda}-\lambda\ln x)^{\frac{1}{\lambda}}&\text{otherwise,}\end{array}\right.$
with $\lambda\in(-\infty,0)$. Considered value $\lambda=-10$.
V Case Studio: Fuzzy implicative association rules
In this section, we apply fuzzy implicative rules to the specific problem of mining association rules. To achieve this, we have implemented an algorithm similar to Apriori [27], incorporating redundancy pruning as described in [28]. This algorithm identifies frequent itemsets based on a minimum fuzzy coverage threshold and then mines association rules by considering minimum thresholds for fuzzy support and confidence. Unlike other approaches, in our framework, support is generally not symmetric. This asymmetry has been taken into account when deriving the final rules. The Python implementation can be found in the following open repository https://github.com/rferper/FIRM, in which we allow the custom selection of fuzzy partitions and operators. Since in Section III-B we have proved that our framework generalizes the crisp and other fuzzy perspectives, instead of comparing with other implementations (see for instance [23, 29]), we will only compare different choices of fuzzy operators. This makes the comparison more robust, since we ensure that the fuzzification process is exactly the same in all evaluations.
For data we have used five datasets from the UCI dataset repository [30]: abalone, iris, wdbc, magic and vehicle. Let us highlight that these datasets have both numeric and categorical variables. For numerical variables we have considered a fuzzy partition with three triangular fuzzy sets defined according the quantiles of the variable and for categorical variables crisp fuzzy partitions, according to the number of classes. For every database we have selected 0.3 as minimum fuzzy coverage and support and 0.8 as minimum fuzzy confidence. The results can be found in Table II.
TABLE II: Results per dataset and pair of operators $(T,I)$. In each cell, it appears the number of fuzzy rules (top) and a tuple with the mean of the four quality measures FCoverage, FSupport, FConfidence and FWRAcc (bottom). Also, under the name of each dataset there is a summary of its specifications: (samples, features).
Dataset $(T,I)$ $(T_{\bm{P}},I_{Y})$ $(T_{\bm{LK}},I_{\bm{LK}})$ $(T_{\bm{P}},I_{\bm{(C_{\lambda}^{H})_{1,0}}})$ $(T_{\lambda}^{\bm{SS}},I^{\bm{SS}}_{k_{\lambda}})$
abalone
(4174,9)
95
(0.36,0.31,0.87,0.18)
96
(0.36,0.33,0.9,0.19)
53
(0.37,0.31,0.83,0.17)
101
(0.36,0.33,0.9,0.19)
iris
(150,4)
14
(0.36,0.32,0.9,0.19)
17
(0.36,0.33,0.9,0.2)
12
(0.35,0.32,0.91,0.2)
16
(0.36,0.32,0.9,0.19)
wdbc
(569,31)
123
(0.37,0.32,0.87,0.17)
164
(0.37,0.33,0.87,0.18)
104
(0.37,0.32,0.88,0.16)
154
(0.37,0.33,0.9,0.18)
magic
(19020,11)
7
(0.37,0.32,0.86,0.17)
13
(0.37,0.33,0.87,0.18)
5
(0.38,0.32,0.85,0.16)
13
(0.37,0.32,0.87,0.18)
vehicle
(846,19)
45
(0.36,0.31,0.87,0.18)
62
(0.36,0.33,0.9,0.19)
29
(0.37,0.31,0.85,0.18)
59
(0.37,0.33,0.9,0.19)
From Table II we can see that the choice of the pair $(T,I)$ affects the total number of rules and the quality metrics. Nonetheless, it is hard to get an idea of how different are the outputs without examining each set of output rules. Also, since all the measures introduced in Section III-C are defined in terms of the pair $(T,I)$, it is not adequate to compare any of these measures when different pairs $(T,I)$ are considered. To overcame this issue we propose a metric for comparing two sets of fuzzy rules using the Jaccard similarity index [31]. Specifically, if $SR_{1}=\{R_{1,1},\dots,R_{1,k}\}$ and $SR_{2}=\{R_{2,1},\dots,R_{2,k}\}$ are two sets of fuzzy rules, we define the percentage of similarity as $SR_{2}=\{R_{2,1},\dots,R_{2,k}\}$ are two sets of fuzzy rules, we define the percentage of similarity as
$Similarity(SR_{1},SR_{2})=\frac{|SR_{1}\cap SR_{2}|}{|SR_{1}\cup SR_{2}|}\cdot 100,$
where we have considered that $R_{1,l}=R_{2,\tilde{l}}$ if they involve the same target class and the same linguistic labels of the feature variables. In Figure 1 we can find the mean of similarity between the set of rules obtained in Table II for the different pair of operators.
Figure 1: Mean similarity between the set of rules obtained in Table II for different choices of $(T,I)$.
From the results in Figure 1 we can derive that the set of output rules obtained by different pairs of operators can be drastically different. Then, apart from modeling the support of a rules as a non-symmetric quantity, these results empirically show that fuzzy implicative rules capture different associations than conjunctive rules even if the same t-norm is used to model the conjunction.
VI Conclusions
In this paper, we have revisited the concept of fuzzy implicative rules, and we have provided a solid theoretical framework for any algorithm interested in mining fuzzy rules modeled as logical conditionals. We have proved that for important characteristics like a monotone support or being the generalization of other frameworks it is necessary to introduce a new property which we have called (MTC). We have studied in depth this new property and we have provided different valid solutions. Although there are plenty of fuzzy implication functions available in the literature, we have exposed that finding meaningful solutions remains a challenging problem, especially if the user wants to model the evaluation of a rule as a non-symmetric quantity. Further, we have developed an open-source Python implementation of our framework for mining fuzzy implicative rules and we have shown that our approach provides distinct knowledge compared to others.
As future work, we want to further study the (MTC) property for obtaining more diverse fuzzy operators that can be considered to model fuzzy implicative rules. Also, we want to take into account fuzzy implicative rules in other rule mining techniques like subgroup discovery or exception rules. More generally, the role of fuzzy implication functions in the study of directional fuzzy rules could be further explored. Finally, we want to consider heuristics to improve the efficiency of mining fuzzy implicative rules.